OpenStudy (anonymous):

Can someone lend me a hand? Algebra 2

OpenStudy (anonymous):

The year is 2033, and intelligent life has been found on Mars! Little green Martians have found their way to Earth and have appointed you as the Ultimate Math Ambassador to their planet. You must help them understand how Earthlings use math

OpenStudy (anonymous):

got to answer 6 questions on this and would like to have a second opinion on my answers :)

OpenStudy (okra77):

ok

OpenStudy (anonymous):

gimme one sec to get the question typed up lol

OpenStudy (anonymous):

Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’ calculations and explain, using complete sentences, what corrections they should make.

OpenStudy (anonymous):

Splott solved it like this: 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 5x + 4 = 14 5x + 4 + 4 = 14 + 4 5x = 18 5 x divided by 5 is equal to 18 divided by 5 x = 18 over 5

OpenStudy (anonymous):

Fizzle solved it like this: 2x + 4 = –3x + 14 2x + 4 – 2x = –3x + 14 – 2x 4 = 14 – 5x 4 – 14 = 14 – 5x – 14 –12 = –5x –12 + 5 = –5x + 5 –7 = x

OpenStudy (anonymous):

were either of them correct?

OpenStudy (anonymous):

didnt splott make an error when he added 3x to the wrong value? 2x + 4 + 3x = –3x + 14 + 3x, =3x should of been added to -3x not 14 right?

ganeshie8 (ganeshie8):

Splott solved it like this: 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x CORRECT 5x + 4 = 14 CORRECT 5x + 4 + 4 = 14 + 4 WRONG, he should have subtracted 4 both sides 5x = 18 5 x divided by 5 is equal to 18 divided by 5 x = 18 over 5

OpenStudy (anonymous):

alright i see it, how about fizzle?

ganeshie8 (ganeshie8):

Fizzle solved it like this: 2x + 4 = –3x + 14 2x + 4 – 2x = –3x + 14 – 2x CORRECT 4 = 14 – 5x CORRECT 4 – 14 = 14 – 5x – 14 CORRECT –12 = –5x WRONG, 4-14 = 10 –12 + 5 = –5x + 5 WEONG, he should gave divided -5 both sides –7 = x

ganeshie8 (ganeshie8):

*have

OpenStudy (anonymous):

Thanks man! i got a couple more questions relating to this :)

OpenStudy (anonymous):

Create your own function to teach the Martians about functions. Your function must contain at least two different operations.

ganeshie8 (ganeshie8):

create something, il give u a hand... fine tuning it :)

OpenStudy (anonymous):

alright ill give it a go!

OpenStudy (anonymous):

5x + 4 = 3x * 2

OpenStudy (anonymous):

will that do?

ganeshie8 (ganeshie8):

no, try something like below :- f(x) = 2x + 2 or f(x) = x^2-3x+2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f(x) = 10x + 6

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

look s good

OpenStudy (anonymous):

Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function.

OpenStudy (anonymous):

so what makes this a legit function? @ganeshie8

ganeshie8 (ganeshie8):

to prove its a legit function, we need to know the definition of function. wats the definition of a function ?

OpenStudy (anonymous):

I think functions will always end up with 1 answer or some rule like that?

OpenStudy (anonymous):

In mathematics, a function[1] is a relation between a set of inputs and a set of permissible outputs with the property that each input is related to exactly one output.

OpenStudy (anonymous):

so everytime we should get the exact same output?

OpenStudy (anonymous):

so would that mean no matter what the x value is the output will be the same?

OpenStudy (anonymous):

or would it mean that there will only be one answer but it may change if the x value is changed?

ganeshie8 (ganeshie8):

the latter one. for every x value, there is a unique y value

ganeshie8 (ganeshie8):

f(x) = 10x + 6 if you put x = 0, u always get, y = 6 u never get any other value

OpenStudy (anonymous):

so everytime you change the x value there will be a y value thats unique to it?

ganeshie8 (ganeshie8):

exactly ! u get the point :)

OpenStudy (anonymous):

alright so how should i explain this in my answer?

ganeshie8 (ganeshie8):

just put wat we discussed in words

OpenStudy (anonymous):

Just say for every x value there is a unique y value?

ganeshie8 (ganeshie8):

u can be more expressive

OpenStudy (anonymous):

and then throw in an example?

ganeshie8 (ganeshie8):

In mathematics, a function[1] is a relation between a set of inputs and a set of permissible outputs with the property that each input is related to exactly one output. that means. for every x value there is a unique y value. there are other relations that are not functions.... u can expa

ganeshie8 (ganeshie8):

the more more words u put, the better

OpenStudy (anonymous):

hey if x = 1 then would y = 16?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yup

ganeshie8 (ganeshie8):

y^2 = x is not a function, why ?

ganeshie8 (ganeshie8):

u may think of it later, after finishing the current assignment :)

OpenStudy (anonymous):

alright i'll give it some thought :)

OpenStudy (anonymous):

Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences.

OpenStudy (anonymous):

so.. f(x) = 10x + 6 y = 10x + 6? /10 y = 6/10?

ganeshie8 (ganeshie8):

nopes

OpenStudy (anonymous):

hmm

ganeshie8 (ganeshie8):

you wanto solve f(3) simply replace x wid 3 in the original function f(x) = 10x + 6 f(3) = 10*3 +6 = 30+6 = 36

OpenStudy (anonymous):

alright I got it. :) and if it was f4 it would of been 46?

OpenStudy (anonymous):

Using complete sentences, describe to the Martians how to find the inverse of your function.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I'm guessing the inverse would like switch everything around?

ganeshie8 (ganeshie8):

you got f(3) = 36 right ? how did u get that ? u simply plugged 3 for x

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

now the inverse is like this : il tell you that 36 is the value of funciton. you're supposed to find wats the input value

ganeshie8 (ganeshie8):

can u tell me for wat value of x, f(x) produces 56 ?

OpenStudy (anonymous):

5

ganeshie8 (ganeshie8):

can u tell me for wat value of x, f(x) produces 106 ?

OpenStudy (anonymous):

10

ganeshie8 (ganeshie8):

now tell me, step by step how did u get 10

OpenStudy (anonymous):

well 10 * 10 = 100 then add 6 to get 106

ganeshie8 (ganeshie8):

you must have mentally done this :- f(x) = 10x+6 106 = 10x + 6 106-6 = 10x 106-6 ------ = x 10

OpenStudy (anonymous):

Yep thats what i did, just couldnt figure out how to put it on paper lol

ganeshie8 (ganeshie8):

finding inverse is the samething just solve for x first

OpenStudy (anonymous):

alright so now i need to solve for 36?

ganeshie8 (ganeshie8):

f(x) = 10x+6 f(x) - 6 = 10x f(x)-6 ------ = x 10

ganeshie8 (ganeshie8):

now, switch f(x) and x

ganeshie8 (ganeshie8):

x-6 ------ = f(x) 10

OpenStudy (anonymous):

36 = 10x + 6 -6 -6 30 = 10x /10 /10 3 = x or x = 3

ganeshie8 (ganeshie8):

this f(x) is ur inverse function

ganeshie8 (ganeshie8):

^ check my work above

OpenStudy (anonymous):

alright so they don't want u to solve it more of just change it into a different equation?

ganeshie8 (ganeshie8):

correct, they want the equation for inverse function

OpenStudy (anonymous):

sorry was afk

ganeshie8 (ganeshie8):

np

OpenStudy (anonymous):

last question

OpenStudy (anonymous):

he Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x).

ganeshie8 (ganeshie8):

wat are our two functions, f(x) , and its inverse g(x) ? can u pull them for me and write them here please ? :)

OpenStudy (anonymous):

yeah np

OpenStudy (anonymous):

f(x) = 10x + 6

OpenStudy (anonymous):

then x-6/10 = 36?

ganeshie8 (ganeshie8):

noo, wahts out inverse functin ?

OpenStudy (anonymous):

whoops sorry lol

ganeshie8 (ganeshie8):

this is our inverse function x-6 ------ = f(x) 10

ganeshie8 (ganeshie8):

sine they're asking us to call the inverse function as g(x), lets call it g(x) x-6 ------ = g(x) 10

ganeshie8 (ganeshie8):

so, our function is f(x) = 10x+ 6 its inverse is g(x) = (x-6)/10

OpenStudy (anonymous):

now we replace x with the inverse in F?

ganeshie8 (ganeshie8):

Next, find f(g(x)) and g(f(x))

OpenStudy (anonymous):

alright

ganeshie8 (ganeshie8):

u should get x for both. give it a try. .... i need to go run a quick errand, brb cya

OpenStudy (anonymous):

alright